In standard form, find the equation of the line passing through (1,8) and
(-2,-1)

Answer by ichudov(499)

You have y=ax+b. You know that for x=1, a*x+b = 8, and for x = -2, a*(-2)+b = -1. It is a standard linear system, for which I have a solver that shows work, on this site.